# Retrieving snow depth distribution by downscaling ERA5 Reanalysis with ICESat-2 laser altimetry

### v1

@@ 4.2 Microscale snow depth variability
Figure 6: Microscale (10 m) snow depth comparison in the west Hardangervidda, April 2008. (a) Transect line, marked white in panels b/c, (b) ALS snow survey validation strip with significant snow depth variations, (c) downscaled snow depth, (d) differences between the ALS data and model output (background: DTM1 DEM). Common horizontal scales of landforms are symbolized by black arrows in panel a.
Figure 6 compares the 2008 ALS snow survey with the downscaled snow depth at a 10-meter resolution. The shown area, located in the western part of flight line 2 (indicated in Figure 1), exhibits distinct microscale landforms, with sheltered depressions hosting thick snow patches (> 8 m) and wind-exposed hilltops featuring thin snow covers (close to 0 m). These terrain features align with the spatial lag that exhibits significant autocorrelation of snow depth in the variogram in Figure 10 (e), corresponding to ca. 100 m and 1.3 km.
The transect line (marked in white) in the downscaled output across this varied terrain visually captures most of the observed pattern very well. Extreme high/low snow depths tend to be biased towards average values, but less so than if scaling calibration is not applied (Figure S.6). While Spearman’s $\rho$ of 0.77 indicates a strong rank correlation, the relatively low $R^{2}$ value of 0.34 and high RMSE of 1.33 m (panel d) suggests a reduced statistical agreement compared to the mesoscale analyses. The KSD of 0.12 for the transect line and a mean deviation of 0.28 m for the difference map indicate that the scaling calibration of the downscaled snow depths, a single function applied to the entire study area and not the shown sub-region specifically, does not fully reproduce extreme values and results in a slight overestimation of average snow depths in this sub-region. The residual differences appear to be correlated with terrain features and match areas with evidence of remaining snow patches in the DTM1, which we detect by comparing the DTM1 with the ALS snow-off data (Figure S.6 e).
@@ 4.3 Temporal variability of snow Depth
The comparison to timeseries focuses on the downscaling model’s ability to propagate information in time rather than space. Note that for this purpose, quantile regression is used in the downscaling model and the uncalibrated model output is shown in Figure 7, to enable analysis of the local bias and its evolution over time. Time series data are shown for the data cell (10 m pixel size) corresponding to the locations of three stations within the Hardangervidda area: Sandhaug (a), Mogen (b), and Haukeliseter (c). The predicted interquartile range Q25–Q75 (IQR, shown in yellow) provides an estimation of the estimated uncertainty of snow depth predictions. Q50 (orange) represents the median prediction. Visually, Q75 shows the best match with the weather station data whereas input ERA5 data over- and Q50 model data underestimate measured snow depths. The $R^{2}$ values of Q50 range from -0.43 (Mogen) to 0.86 (Haukeliseter), thus don’t describe model performance well. Bias is primarily observed for high snow depth values during the peak snow season whereas the snow-free season is captured accurately. The peak snow bias is different for the three sites but has a consistent magnitude over time at each site and there are no indications of a better fit for the period with training samples (late 2018 to late 2022) compared to earlier years, indicating that the bias could be reduced or removed entirely by a scaling calibration such as applied to the result maps above. While the (uncalibrated) model exhibits bias in predicting the absolute snow depths, the high Spearman’s $\rho$ (0.91 to 0.94) across all three stations suggests that the downscaling successfully maintains the relative ordering of snow depth time series.
Figure 7: Time-series validation of snow depth for three weather stations Sandhaug (a), Mogen (b), and Haukeliseter (c). Blue: measured monthly mean and minimum/maximum values (light blue), orange: median downscaled snow depth (Q50) and interquartile range (IQR) for 25th–75th quantiles (yellow). Note: Haukeliseter and Sandhaug data availability starts from 2015 and 2014, respectively, and some years have incorrect measurements.
@@ 4.4 Validating snow occurrence
Figure 8 provides a visual comparison between the model output and Sentinel-2 imagery for June 2020 in the lake Møsvatn area (Figure 1), showing the model’s accuracy in predicting the occurrence of snow during the rapid melt period. The downscaled snow map from our model visually shows a high level of agreement with the satellite snow extent on June 24, 2020, capturing snow patches that align with topographical features (Figure 8 c,d). Panel (e) and (f) show that east-facing slopes retain more snow compared to west-facing slopes, a pattern that our model successfully captures. Closer inspection reveals minor discrepancies in snow distribution. Some of these differences can be explained by the binary nature of Sentinel-2 snow cover and a temporal mismatch, with the Sentinel-2 image taken on a specific day and the model output representing a monthly mean snow depth for June. In particular, the model tends to systematically overestimate snow coverage on specific terrain, such as steep slopes (indicated by the black arrow), and underestimate it in flatter areas (indicated by the white arrow). For a dynamic view of snow occurrence variability, please refer to the video in the supplementary materials.
Figure 8: Validation of snow occurrence using Sentinel-2 imagery for the Lake Møsvatn area. (a) modelled snow cover distribution for June 2020, (b) Sentinel-2 image on June 24th, 2020. Panels (c) and (d) show local discrepancies in snow distribution, (e) and (f) show correctly modelled snow retention on east-facing versus west-facing slopes. The overestimation on steep slopes and underestimation in flatter terrain are annotated with black and white circles, respectively.
@@ 4.5 SHAP analysis and variogram assessment
The most informative relationships between the input features and the predicted subgrid variability are listed in descending order. Elevation (h_te_best_fit) emerges as the most significant factor, with higher elevations positively influencing subgrid snow depth (Figure 9 a). Snow depth from ERA5 Land also has a high impact, however, both very thin snow depth and thick localized snow depth exist (Figure 9 a,b) where (sde_era) is high, owing to the low spatial resolution of the data that does not capture topography. Notably, relatively thicker snow (indicated by high SHAP values in Figure 9 b,c) is estimated by the downscaling model in concave terrain (negative TPI) with an east-facing slope. In contrast, thinner snow (lower SHAP values) is associated with convex (positive TPI) and west-facing slopes. The model estimates higher snow depths for the two northernmost ALS flight strips of the study area than the southern part (Figures 9 d, 5). The positive cumulative wind aspect factor (wuf_positive) contributes to preferential snow accumulation and shows a stronger influence than the negative wind aspect factor (wf_negative, Figure 9 f,g). Most wind aspect factor data points have small values indicative of low wind speeds. These show much variability in their SHAP value, indicating unexplainable variance. SHAP cannot distinguish the contributions from correlated features such as the apparent importance and positive influence of slope, which might partly be caused by correlated features such as elevation and curvature. The phenomenon of having less snow on very steep slopes (e.g. >50°) is not observed (Figure 9 a), rather, the model associates high slopes with a positive SHAP value that indicates greater snow depths.
@@
Figure 10: Spatial correlation from semivariograms based on six ALS validation strips (2008) and downscaling model outputs at 10 m resolution. The semivariograms are fitted to the sum of the spherical model and Gaussian model from short to long ranges. Note the limited capture of variance in the transverse direction due to the ALS strips’ constrained width.
@@
The empirical semivariogram in Figure 10 shows that the downscaling models based on all DEM versions adequately capture the spatial variability and autocorrelation of snow depths in the Hardangervidda area, but suggests stronger spatial dependence at short ranges for ALS data (79% variance correlated in 95 m lag) than downscaled snow depths (ca. 60% of the variance in 108 m lag). The distance of ca.  100 m corresponds to the wavelength of depression features commonly observed in the area (Figure 6). Overall, the downscaling model exhibits 10% - 13% less variance than ALS data, suggesting that the models perform somewhat worse in representing the true snow depth spatial variability. Hereby, the Norwegian national DEMs perform better (3% higher semi-variance values) than the global DEMs. At a distance of ca. 1,300 m, the semi-variance for all datasets approaches the sill. This distance aligns closely with the typical size of hillslopes in this region, as depicted in Figure 6 and Figure 5. The interpolated SeNorge product, based on precipitation gradients, temperature–elevation lapse-rates and disregarding wind processes, has a too coarse resolution to resolve hillslope-scale snow depth variability. It can explain 69% of the total variance occurs over a spatial lag of 16,000 meters. This distance corresponds to the maximum local relief from the mountaintops to the lake surface (Figure 5).
@@ 5.1 Bias in elevation datasets
This study shows that ICESat2 ATL08 data is not just capable to retrieve snow depths, thus confirming previous findings by other authors [13, 16, 82, 3], but it can be used to co-register and rectify DEMs. Previous studies have in common that retrieved snow depths are biased by remaining discrepancies and errors in the DEM/ICESat-2 data such as a slope-dependent bias. The bias correction introduced in our workflow can remove errors that depend on first and second-order derivatives (slope/curvature). Several studies pointed to ICESat-2 as the primary source of this bias, as ICESat-2 ATL08 data tends to underestimate surface height under certain conditions. Moudrý et al. [61] attributed the error to the presence of clouds and corresponding increasing atmospheric scattering effects leading to an increased photon travel time and, consequently, underestimating terrain height. Furthermore, the 100 m segment length of ATL08 (e.g. h_te_mean used by Enderlin et al. 16) is considered insufficient for accurately mapping steep and rugged terrain [3]. Therefore, Enderlin et al. [16] recommended using ATL08 only in areas with relatively low slopes and sparse vegetation cover. We also observed that bias is higher in the Hardangervidda area than in other, flatter parts of Norway, such as Finnmark (not shown in this paper). In our data sample, a large number of negative snow depth measurements were present before bias correction. Directly excluding these negative snow depths (without further bias correction) would result in unbalanced training samples. A correction solely based on the slope, as proposed by Enderlin et al. [16] and Tian and Shan [82], did not remove bias sufficiently for the ICESat-2 measurements to be used as model input. We find resolution-dependent biasing effects for the tested DEMs with regard to curvature that are essential to correct for accurate snow depth retrieval (Supplementary Section B). For Norway, we found that the spatial resolution of the ATL08 data is better than COP30 (30 m) when using h_te_best_fit elevation values that are fitted to the middle 20 m of the 100-m-ATL08 segment (Supplementary Section B).
@@
Also the snow-off DEMs hold potential limitations for ICESat-2-derived snow depths in certain terrain, especially high mountain or Arctic areas where the snow-free season is short and some snow patches remain during summer. Typically, some snow at the highest elevations is of little concern for general-purpose DEM data distributors who need to balance requirements of numerous DEM applications. The DTM1 was acquired during June/July according to metadata, under summer conditions. From comparison with ALS snow-off data (Figure S.6), we find that the Hardangervidda area was not entirely snow-free for any of the four tested DEMs nor the Arctic DEM [73]. Remaining snow on the supposedly snow-free reference results in an underestimation of snow depth in parts of these areas, with typical topographic signatures (e.g., shaded depressions). Additionally, in areas with long snow cover duration there are fewer snow-free ICESat-2 segments available. Consequently, the bias correction introduced in this study may not be able to correctly capture and remove this bias, leading to underestimation of high snow depths. When comparing our downscaled snow depths with the ALS validation data, we find that on the microscale, the residuals are highly associated with a preferential snow deposition pattern. The systematic underestimation of high snow depths is even more prominent in the downscaled output if the calibration step is not applied (Figures S.8, S.9). The bias correction and calibration steps only indirectly correct for the biasing effect of remaining snow patches on retrieved/downscaled snow depths, and they also address other biasing effects (see section 5.2). Future work should thus include a targeted correction of the DEM data for remaining snow patches. This would improve our results and potentially reduce the need for the currently included calibration step.
Enderlin et al. [16] deemed COP30 data too imprecise for estimating snow depth. Encouragingly, our results demonstrate that the COP30 and FAB DEMs improved after bias correction, and perform comparably with DTM1 and DTM10 at mesoscale in treeless areas (Figure S.7). We attribute this to the bias correction step in our workflow which successfully removes bias.
@@
This study only validates the downscaled snow depths above the tree line and further work is needed to assess the performance of the proposed workflow in forested areas where ALS validation data is available. We expect that more, or different predictors may be needed to capture information on snow–canopy processes [53] influencing the accumulation and ablation of snow. High-resolution lidar data allows for the removal of vegetation, but global-scale DEMs are usually based on photogrammetry or radar interferometry, which may not exclude vegetation cover. There are efforts to create global-scale vegetation-free DEMs like the FAB DEM included in this study, and further developments are expected from machine-learning algorithms. Currently, for the high elevation accuracy required for snow depth retrieval, we recommend to be cautious about using COP30 and FAB in forested areas, as vegetation is hardly (fully) removed.

### v2

@@ 4.2 Microscale snow depth variability
Figure 6: Microscale (10 m) snow depth comparison in the west Hardangervidda, April 2008. The validated area is located along flight line b, as shown in the study area map in Figure 1 and mesoscale snow depth profile in Figure 5. (a) Transect line, marked white in panels b/c, (b) ALS snow survey validation strip with significant snow depth variations, (c) downscaled snow depth, (d) differences between the ALS data and model output (background: DTM1 DEM). Common horizontal scales of landforms (95m and 1.3 km) are symbolized by black arrows in panel a.
Figure 6 compares the 2008 ALS snow survey with the downscaled snow depth at a 10-meter resolution. The area shown, located in the western part of flight line 2 (indicated in Figure 1), exhibits distinct microscale landforms, with sheltered depressions hosting thick snow patches (> 8 m) and wind-exposed hilltops featuring thin snow covers (close to 0 m). This area was also shown in Melvold and Skaugen [60, Figure 3] to address the effect of spatial resolution on snow depth representation. The terrain features align with the spatial lag that exhibits significant autocorrelation of snow depth in the variogram in Figure 10 (e), corresponding to ca. 100 m and 1.3 km.
The transect line (marked in white) in the downscaled output across this varied terrain visually captures most of the observed pattern. Extreme deep/shallow snow depths tend to be biased towards average values, but less so than if scaling calibration is not applied (Figure S.6). While Spearman’s $\rho$ of 0.77 indicates a strong rank correlation, the relatively low $R2$ value of 0.34 and high RMSE of 1.33 m (panel d) suggests a reduced statistical agreement compared to the mesoscale analyses. The KSD of 0.12 for the transect line and a mean deviation of 0.28 m for the difference map indicate that the scaling calibration of the downscaled snow depths, a single function applied to the entire study area and not the shown sub-region specifically, does not fully reproduce extreme values and results in a slight overestimation of average snow depths in this sub-region. The residual differences appear to be correlated with terrain features and match areas with remaining snow patches in the DTM1, which we detect by comparing the DTM1 with the ALS snow-off data (Figure S.6 e).
@@ 4.3 Temporal variability of snow depth
The downscaling model is also able to propagate information in time rather than only space. For this purpose, we used quantile regression for the downscaling model. The uncalibrated model output is shown in Figure 7, to visualize the local bias and its evolution over time. Time series data are shown for the data cell (10 m pixel size) corresponding to the locations of the three stations within the Hardangervidda area: Sandhaug (a), Mogen (b), and Haukeliseter (c). The predicted interquartile range Q25–Q75 (IQR, shown in yellow) provides an estimate of the snow depth predictions uncertainty. Q50 (blue line) represents the median prediction. Visually, Q75 shows the best match with the weather station data whereas ERA5 data over- and Q50 model data underestimate measured snow depths (Table 2). ($R2$ of 0.71, 0.60, 0.76) The downscaled model (both Q50 and Q75) generally performs similarly to the original ERA5 Land data across all stations. $R2$ values range from -0.43 (Q50, Mogen) to 0.86 (Q50, Haukeliseter), and do not describe performance well due to significant biases at the individual site level. While the (uncalibrated) model exhibits bias in predicting the absolute snow depths, the high Spearman’s $\rho$ (0.91 to 0.95, ERA5: 0.92 to 0.96) across all three stations suggests that the downscaling successfully maintains the relative ordering of snow depth time series. The better agreement of Q75 with observations (compared to Q50) could reflect either (1) a systematic  25% underestimation bias in the model, or (2) the local station measurements corresponding to the Q75 percentile of the model’s predictions for similar conditions. Bias is primarily observed for high snow depth values during the peak snow season whereas the snow-free season is captured accurately. The peak snow bias is different for the three sites but has a consistent magnitude over time at each site. There are no indications of a better fit for the training period (late 2018 to late 2022) compared to earlier years. This suggests that the bias could be reduced or removed entirely by a scaling calibration such as applied to the maps resulting from the spatial propagation presented above.
Figure 7: Time-series validation of snow depth for three weather stations Sandhaug (a), Mogen (b), and Haukeliseter (c). Blue Dot: measured monthly mean and minimum/maximum values (light blue), Blue line: median downscaled snow depth (Q50) and interquartile range (IQR) for 25th–75th quantiles (yellow). Note: Haukeliseter and Sandhaug data availability starts from 2015 and 2014, respectively, and some years have missing data and incorrect measurements.
Table 2: Statistical comparison of monthly snow depth estimates from different methods (Downscaled output and ERA5 Land) against weather station snow depth. RMSE is in metres.
Model — Q50 Model — Q75 ERA5 Land
Station R2 KSD $\rho$ RMSE R2 KSD $\rho$ RMSE R2 KSD $\rho$ RMSE
Sandhaug 0.29 0.20 0.94 0.39 0.71 0.17 0.95 0.29 0.73 0.21 0.96 0.44
Mogen -0.43 0.22 0.91 0.28 0.60 0.19 0.91 0.23 0.47 0.31 0.92 0.46
Haukeliseter 0.86 0.10 0.94 0.26 0.76 0.23 0.94 0.38 0.37 0.36 0.94 0.79
@@ 4.4 Validating snow occurrence
Figure 8 provides a visual comparison between the downscaling model output and Sentinel-2 imagery for June 2020 in the lake Møsvatn area (Figure 1). The figure shows the model’s accuracy in predicting the occurrence of snow during the rapid melt period. The downscaled snow map from our model visually shows a high level of agreement with the satellite snow extent on June 24, 2020, capturing the remaining snow patches aligning with topographical features (Figure 8 c,d). Panel (e) and (f) show that east-facing slopes retain more snow compared to west-facing slopes, a pattern that our model successfully captures. Closer inspection reveals that the model retains a thin snow layer in most areas as well as minor discrepancies in snow distribution. Some of these differences can be explained by the binary nature of Sentinel-2 snow cover and a temporal mismatch, with the Sentinel-2 image taken on a specific day and the model output representing a monthly mean snow depth for June. In particular, the model tends to systematically overestimate snow coverage for specific terrain (indicated by the white circle), such as steep slopes, and underestimates snow coverage for certain terrain features (white circle). To express the match quantitatively, we converted the model output to snow occurrence/probability using a threshold of 0.3 m for full snow cover, and the Sentinel-2 data to a binary reference by computing the Normalised Snow Difference Index (using a threshold of 0.42, as recommended by the Sentinel-2 data provider Copernicus). For the area shown in Figure  8 (extent of panels a,b), we receive a Matthews Correlation Coefficient of 0.35 and an Area Under the Curve (AUC) value of 0.76, both indicating acceptable model performance. For a dynamic view of snow occurrence variability, please refer to the video in the supplementary materials.
Figure 8: Validation of snow occurrence using Sentinel-2 imagery for the Lake Møsvatn area (annotated in Figure 1). (a) modelled snow cover distribution for June 2020, (b) Sentinel-2 image on June 24th, 2020. Panels (c) and (d) show local discrepancies in snow distribution, (e) and (f) show correctly modelled snow retention on east-facing versus west-facing slopes. The overestimation on steep slopes and underestimation in flatter terrain are annotated with white circles.
@@ 4.5 SHAP analysis and variogram assessment
The most informative relationships between input features and predicted subgrid variability are listed in descending order. Elevation (h_te_best_fit) emerges as the most significant factor, with higher elevations positively influencing subgrid snow depth (Figure 9 a). Snow depth from ERA5 Land also has a high impact, however, both very shallow snow depth and deep localized snow depth exist (Figure 9 a,b) where (sde_era) is high, owing to the low spatial resolution of the data not capturing the variable topography. Notably, relatively thicker snow (indicated by high SHAP values in Figure 9 b,c) is estimated by the downscaling model in concave terrain (negative TPI) with an east-facing slope. In contrast, thinner snow (lower SHAP values) is associated with convex (positive TPI) and west-facing slopes. The model estimates deeper snow depths for the two northernmost ALS flight strips of the study area than the southern part (Figures 9 d, 5). The positive cumulative wind aspect factor (wuf_positive) contributes to preferential snow accumulation and shows a stronger influence than the negative wind aspect factor (wf_negative, Figure 9 f,g). Most wind aspect factor data points have small values indicative of low wind speeds. These show much variability in their SHAP value, indicating unexplainable variance. The SHAP method cannot distinguish contributions from correlated features such as the apparent importance and positive influence of slope, which might partly be caused by correlated features such as elevation and curvature. The phenomenon of having less snow on very steep slopes (e.g. >50°) is not observed (Figure 9 a), rather, the model associates high slopes with a positive SHAP value that indicates greater snow depths.
@@
Figure 10: Spatial correlation from semivariograms based on six ALS validation strips (2008) and downscaling model outputs at 10 m resolution. Note the limited capture of variance in the transverse direction due to the ALS strips’ constrained width (500m). The vertical bar on crosses represents one standard deviation where data are sufficient to compute the statistic.
@@
The empirical semivariogram in Figure 10 shows that the downscaling models based on all DEM versions adequately capture the spatial variability and autocorrelation of snow depths in the Hardangervidda area, but suggests stronger spatial dependence at short ranges for ALS data than downscaled snow depths (ca. 80% and 50% of the total variance at 100 m lag). The distance of ca.  100 m corresponds to the wavelength of depression features commonly observed in the area (Figure 6). Overall, the downscaling model exhibits ca. 20% less variance than the ALS data. At a distance of ca. 1,300 m, the semi-variance for all datasets approaches the sill. This distance aligns closely with the typical size of hillslopes in this region, as depicted in Figure 6. The interpolated SeNorge product, based on precipitation gradients, temperature–elevation lapse-rates and disregarding wind processes, has a resolution too coarse to resolve the hillslope-scale snow depth variability.
@@ 5.1 Bias in elevation datasets
This study shows that ICESat2 ATL08 data is capable to retrieve snow depths, thus confirming previous findings by other authors [15, 18, 92, 4], while also improving snow depth retreival through co-registeration and rectification of DEMs. Previous studies have in common that retrieved snow depths are biased by remaining discrepancies and errors in the DEM/ICESat-2 data such as a slope-dependent bias. The bias correction introduced in our workflow can reduce errors that depend on first and second-order derivatives (slope/curvature). Several studies pointed to ICESat-2 as the primary source of this bias, as ICESat-2 ATL08 data tends to underestimate surface height under certain conditions. Moudrý et al. [67] attributed the error to the presence of clouds and corresponding increasing atmospheric scattering effects leading to an increased photon travel time and, consequently, underestimating terrain height. Furthermore, the 100 m segment length of ATL08 (e.g. h_te_mean used by Enderlin et al. 18) is considered insufficient for accurately mapping steep and rugged terrain [4]. Therefore, Enderlin et al. [18] recommended using ATL08 only in areas with relatively low slopes and sparse vegetation cover. We also observed that bias is higher in the Hardangervidda area than in other, flatter parts of Norway, such as Finnmark (not shown in this paper). In our data sample, a large number of negative snow depth measurements were present before bias correction. Directly excluding these negative snow depths (without further bias correction) would result in unbalanced training samples. A correction solely based on the slope, as proposed by Enderlin et al. [18] and Tian and Shan [92], did not remove bias sufficiently for the ICESat-2 measurements to be used as model input for the downscaling. We find resolution-dependent biasing effects for the tested DEMs with regard to curvature that are essential to correct for accurate snow depth retrieval (Supplementary Section B). For Norway, we found that the spatial resolution of the ATL08 data is better than COP30 (30 m) when using h_te_best_fit elevation values fitted to the middle 20 m of the 100-m-ATL08 segment (Supplementary Section B).
@@
Also the snow-off DEMs hold potential limitations for ICESat-2-derived snow depths in certain terrain, especially high mountain or Arctic areas where the snow-free season is short and some snow patches remain during summer. Typically, some snow at the highest elevations is of little concern for general-purpose DEM data distributors who need to balance requirements of numerous DEM applications. The DTM1 was acquired during June/July according to metadata, under summer conditions. From comparison with ALS snow-off data (Figure S.6), we find that the Hardangervidda area was not entirely snow-free for any of the four tested DEMs nor the Arctic DEM [82]. Remaining snow on the supposedly snow-free reference results in an underestimation of snow depth in parts of these areas, with typical topographic signatures (e.g., shaded depressions). Additionally, in areas with long snow cover duration there are fewer snow-free ICESat-2 segments available. Consequently, the bias correction introduced in this study may not be able to correctly capture and remove this bias, leading to underestimation of high snow depths. When comparing our downscaled snow depths with the ALS validation data, we find that on the microscale, the residuals are highly associated with a preferential snow deposition pattern. The systematic underestimation of deep snow depths is even more prominent in the downscaled output if the calibration step is not applied (Figures S.8, S.9). The bias correction and calibration steps only indirectly correct for the bias effect of remaining snow patches on retrieved/downscaled snow depths. They also address other bias effects (see section 5.2). Future work should thus include a targeted correction of the DEM data for remaining snow patches. This would improve our results and potentially reduce the need for the current calibration step.
Enderlin et al. [18] deemed COP30 data too imprecise for estimating snow depth. Encouragingly, our results demonstrate that the COP30 and FAB DEMs improved after bias correction, and performed comparably with DTM1 and DTM10 at mesoscale in treeless areas (Figure S.7). We attribute this to the bias correction step in our workflow which successfully removes bias for all four tested DEMs (Supplementary material). We are thus confident that the regression-based bias correction is transferable to other regions for the same (global COP30 and FAB DEMs) or different DEM products. As the training area and sample were very large and diverse in this study (entire Norway), further research on the bias-correction method could examine the applicability to different landscapes/topography or the influence of the size of the area and the ICESat-2 snow-off sample. We note that the remaining elevation error of 0.48–0.62 m (NMAD after bias correction) for DTM1 and FAB, respectively, may be acceptable for deep snow or a snowpack highly affected by wind redistribution (e.g. this study), but can exceed the thickness of shallow snow packs. The influence of uncertain individual samples is mitigated by the approach of this study, where the ICESat-2 measurements are used as a training sample for a downscaling algorithm rather than on their own. The uncertainty may be reduced for regions with less snow depth variability, lowering de-facto the absolute magnitude of discrepancies between observations and model outputs. Further research is needed to better understand the effect of the snow depth sample uncertainty on the downscaled snow depth maps, in particular the poor representation of extremes. It should also be noted that in this study, terrain parameters used for downscaling are based on the DTM10. The spatial resolution of the COP30 and FAB DEMs is coarser (30m) than our model output resolution (10m) and they were used for snow depth retrieval only, not for downscaling. When applying our workflow to other regions the spatial resolution of the output snow depth maps will therefore be limited by the available DEM.
@@
This study only validates the downscaled snow depths above the tree line and further work is needed to assess the performance of the proposed workflow in forested areas where ALS validation data is available. We expect that more, or different predictors may be needed to capture information on snow-canopy processes [58] influencing the accumulation and ablation of snow. High-resolution lidar data allows for the removal of vegetation, but global-scale DEMs are usually based on photogrammetry or radar interferometry, which may not exclude vegetation cover. There are efforts to create global-scale vegetation-free DEMs like the FAB DEM included in this study, and further developments are expected from machine-learning algorithms. Currently, for the high elevation accuracy required for snow depth retrieval, we recommend to be cautious about using COP30 and FAB in forested areas, as vegetation is hardly (fully) removed.